OpenStudy (marsmcnugget):

A porch has the shape of a quarter circle with a diameter of 5 ft. What is the approximate area of the porch, rounded to the nearest square foot? Use 3.14 to approximate

OpenStudy (jhannybean):

Area of a circle : \(A = \pi r^2\) So the area of a quarter of a circle would be: \(A = \dfrac{1}{4} \pi r^2\)

OpenStudy (jhannybean):

\[A = \pi r^2 = \pi D\]